# Hierarchical Gaussian Mixture Normalizing Flow Modeling for Unified Anomaly Detection

### v1

@@ 0.A.1 Necessities and Motivations
Learning intra-class mixed class centers. Finally, we consider that in real-world scenarios, even one object class may contain diverse normal patterns. We think modeling intra-class distribution by mixture Gaussian prior should also be beneficial for the results (see Tab. 3a). Moreover, as we further explain in Sec. 0.A.3, another consideration is to guarantee the effectiveness of anomaly determination. The inter-class Gaussian mixture modeling can’t effectively guarantee the anomalies that fall into the inter-class Gaussian mixture distribution to be correctly recognized. To this end, we further model the intra-class Gaussian mixture distribution for each class to ensure that the normal distribution of each class still remains compact. Therefore, even if anomalies fall into the inter-class Gaussian mixture distribution, they are usually in the low-density regions among the inter-class class centers. So, we can still ensure that anomalies are out-of-distribution through intra-class log-likelihoods.
@@ 0.A.2 Supervision Information
0.A.3 The Way to Guarantee Anomalies Out-of-Distribution.
@@ 0.A.5 Model Complexity
0.A.6 Real-world Applications
@@ Appendix 0.G Additional Results
Table 6: Anomaly detection and localization results on MVTecAD. All methods are evaluated under the one-for-one setting. /normal-⋅normal-$\cdot$/$\cdot$ / ⋅ means the image-level and pixel-level AUROCs.

### v2

@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Natural Science Fund of China (62371295), the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and the Science and Technology Commission of Shanghai Municipality (22DZ2229005).
@@ 0.A.1 Necessities and Motivations
Learning intra-class mixed class centers. Finally, we consider that in real-world scenarios, even one object class may contain diverse normal patterns. We think modeling intra-class distribution by mixture Gaussian prior should also be beneficial for the results (see Tab. 3a). Moreover, as we further explain in Sec. 0.A.4, another consideration is to guarantee the effectiveness of anomaly determination. The inter-class Gaussian mixture modeling can’t effectively guarantee the anomalies that fall into the inter-class Gaussian mixture distribution to be correctly recognized. To this end, we further model the intra-class Gaussian mixture distribution for each class to ensure that the normal distribution of each class still remains compact. Therefore, even if anomalies fall into the inter-class Gaussian mixture distribution, they are usually in the low-density regions among the inter-class class centers. So, we can still ensure that anomalies are out-of-distribution through intra-class log-likelihoods.
@@ 0.A.2 Supervision Information
0.A.3 More Discussions with “identical shortcut”
The “identical shortcut” is essentially caused by the leakage of abnormal information. The process of reconstruction is to remove abnormal information in the input, resulting in the failure of reconstruction in abnormal regions. But if the abnormal features are leaked into the output, this will result in the reconstruction network directly returning a copy of the input as output. This issue usually can be addressed by masking, such as the neighbor masking mechanism in UniAD [42]. However, the “homogeneous mapping” is a specific issue in normalizing flow (NF) based AD methods. In previous NF-based AD methods, the latent feature space is uni-modal. When used for unified anomaly detection, we need to map different class features to the single latent center, this may cause the model more prone to take a bias to map different input features to similar latent features. Thus, with the bias, the log-likelihoods of abnormal features will become closer to the log-likelihoods of normal features, causing normal misdetection or abnormal missing detection. We call this phenomenon as the “homogeneous mapping” issue, rather than casually introducing it. Moreover, as analyzed in Sec. 3.2, we provide a reasonable explanation from the perspective of the formula in normalizing flow. To address this issue, we propose the hierarchical Gaussian Mixture modeling approach, the key designs in our method are completely different from those in UniAD. As the causes and solutions of the two issues are significantly different, “homogeneous mapping” is not intrinsically equal to “identical shortcut”.
0.A.4 The Way to Guarantee Anomalies Out-of-Distribution.
@@ 0.A.6 Model Complexity
0.A.7 Real-world Applications
@@ Appendix 0.G Additional Results
Table 6: Anomaly detection and localization results on MVTecAD. All methods are evaluated under the one-for-one setting. $\cdot$/$\cdot$ / ⋅ means the image-level and pixel-level AUROCs.